OpenStudy (anonymous):

I need a step by step walk through on how to solve w^2-18w^2-2=0. The question ask was is the solutions of w

OpenStudy (anonymous):

\[ w^2-18w^2-2=0\] combine like terms \[-17w^2-2 = 0\] \[w^2 = \frac{2}{-17}\] \[w = -i\sqrt{\frac{2}{17}}\] and \[w = i\sqrt{\frac{2}{17}}\]